OpenStudy (anonymous):

is anyone on here i have an easy question in matlab about defining a function

OpenStudy (anonymous):

function [ f ] = f( x ) f(x) = (x^6)-(11.*x^3)+(17.*x^2)-(7.*x)+1; end

OpenStudy (anonymous):

error not enough input arguments when trying to call f in matlab

OpenStudy (anonymous):

here is the program

OpenStudy (anonymous):

function y=gss(f,a,b,k) g=(sqrt(5)-1)/2; x1 = a+(1-g)*(b-a); x2 = a+g*(b-a); f1=f(x1);f2=f(x2); for i=1:k if f1<f2 b=x2; x2=x1; x1=a+(1-g)*(b-a); f2=f1; f1=f(x1); else a=x1; x1=x2; x2=a+g*(b-a); f1=f2; f2=f(x2); end end y=(a+b)/2;